OpenStudy (anonymous):

A spherical ball of radius 6 cm is to be divided into eight equal parts by cutting it along the axis as shown. Find the surface area of each piece in cm2.

OpenStudy (anonymous):

surface area of sphere=4pi r^2 =4*36 r^2=144 r^2

OpenStudy (anonymous):



OpenStudy (anonymous):

so here are 8 divisions=144pie/8=18pie.but it is not correct

OpenStudy (anonymous):

You are only including the area of the round part.

OpenStudy (anonymous):